OpenStudy (anonymous):

If two ladders are placed end to end, their combined height is 46.5 feet. One ladder is 3.5 feet shorter than the other ladder. What are the heights of the two ladders?

OpenStudy (anonymous):

Total height of ladder = 46.5 feet approximately height of one ladder = 46.5 divide by 2 =23.25 feet one ladder is 3.5 feer shorter,= 23.25 -3.5 = 19.75 feet so another ladder should be , = 23.25+3.5 =26.75 feet